OpenStudy (anonymous):

3(x-7)+5=-2x-8... help?

OpenStudy (anonymous):

3x -21 +5 = -2x -8 3x -16 = -2x -8 5x = 8 x = 8/5

OpenStudy (aroub):

3x-21+5=-2x-8 3x+2x=21-5-8 5x=8 x=8/5

OpenStudy (anonymous):

thanks<3